# Strada-LLM: Graph LLM for traffic prediction

### v1

@@ IV-C Baselines
  • STSGCN  [22]: Spatial-Temporal Synchronous Graph Convolutional Layers are used to process the traffic as a spatio-temporal tensor.
@@
  • GMAN  [23]: The gating fusion mechanism redefines the spatio-temporal attention block.
  • MTGNN  [65]: Alternating use of graph convolution and temporal convolution modules.
@@
  • GTS  [25]: The underlying graph structure is learned among multiple time series, and the corresponding time series is simultaneously predicted with DCRNN [1].
@@
  • STEP  [66]: Adopts the graph wavenet model [10], integrated into the transformer backbone and pre-training scheme.
@@
  • FlashST [14]: The data distribution shift is learned by leveraging prompt-tuning.
@@
To compare with the baselines, we adopt MAE, RMSE, and MAPE prediction error [14].
IV-D LLM Model Results
The long-term traffic prediction accuracy was evaluated across four datasets: PeMS-Bay, METR-LA, PEMS07(M), and Brussels. The last dataset is private. Strada-LLM, as an LLM should be benchmarked jointly for all datasets. So, we report the performance in two manners:
  1. 1.
    knowledge adaptation: Like every LLM, we pre-trained Strada-LLM on the datasets PeMS03, PeMS04, PeMS07, and PeMS08 then fine tuned on the PeMS-Bay, METR-LA, PEMS07(M), and Brussels datasets in few-shot fashion; we refer to this approach as LLM setup.
  2. 2.
    fully-supervised: We trained our proposed model on a single dataset in a fully-supervised fashion (referred to as $Strada-LLM\left[Solo\right]$ in this section).
The prediction results for the datasets METR-LA and PEMS-Bay are demonstrated in Tables III and IV, correspondingly.
The ability to perform long-term forecasting and being robust to new domains are important qualities for an LLM. As shown in Tables III and IV, the proposed LLM has competitive performance compared to Step [66], which is the corresponding transformer-based fully-supervised method. On the other hand, compared to the best prompt-tuning method FlashST [14], taking the global structure of the graph into account brings forth a definite advantage. Moreover, although Strada-LLM and STEP [66] are both transformer-based models, the introduction of normalization blocks like RMSNorm [57] and Rotary Positional Encoding (RoPE) differentiates them. Additionally, while STEP [66] is an encoder-decoder model, Strada-LLM is a decoder-only model, which makes it optimized in terms of performance for inference [67]. As we can see in Tables IIIIV, the Strada-LLM manages to outperform the baseline models on the aforementioned datasets. Specifically, on the PeMS-Bay dataset, our proposed model achieves an RMSE of 3.94 for 1-hour prediction, corresponding to an improvement of approximately $16\%$ compared to the prompt-tuning baseline [14].
TABLE III: Traffic forecasting comparison with the state-of-the-art methods on the METR-LA dataset
Baselines Training Time(H)a 15 min 30 min 60 min
@@ IV-D LLM Model Results
Supervised STSGCN [22] - 3.31 7.62 8.06 4.13 9.77 10.29 5.06 11.66 12.91
GMAN [23] - 2.80 5.55 7.41 3.12 6.49 8.73 3.44 7.35 10.07
MTGNN [65] 3 2.69 5.18 6.88 3.05 6.17 8.19 3.49 7.23 9.87
GTS [25] 6 2.67 5.27 7.21 3.04 6.25 8.41 3.46 7.31 9.98
STEP[66] 18 2.61 4.98 6.6 2.96 5.97 7.96 3.37 6.99 9.61
LLM FlashST [14] 0.5 2.84 5.57 7.45 3.19 6.43 9.04 3.60 7.44 10.68
Strada-LLM(ours) 0.5 2.73 4.95 6.54 2.99 5.84 7.87 3.43 6.73 9.47
The reported numbers were reproduced on our hardware. For the LLM-based models, the entire duration of the few-shot learning process is reported.
TABLE IV: Traffic predictions comparison with the state-of-the-art methods on PEMS-Bay dataset
Baselines Training Time(H)a 15 min 30 min 60 min
[…]
Supervised STSGCN [22] - 1.44 3.01 3.04 1.83 4.18 4.17 2.26 5.21 5.40
GMAN [23] - 1.34 2.91 2.86 1.63 3.76 3.68 1.86 4.32 4.37
MTGNN [65] 3 1.32 2.79 2.77 1.65 3.74 3.69 1.94 4.49 4.53
GTS [25] 6 1.34 2.83 2.82 1.66 3.78 3.77 1.95 4.43 4.58
STEP[66] 18 1.26 2.73 2.59 1.55 3.58 3.43 1.79 4.20 4.18
LLM FlashST [14] 0.5 1.37 2.96 2.88 1.72 3.98 3.89 2.02 4.69 4.79
Strada-LLM (Ours) 0.5 1.28 2.77 2.55 1.63 3.14 3.27 1.83 3.94 4.09
The reported numbers were reproduced on our hardware. For the LLM-based models, the entire duration of the few-shot learning process is reported.
[…]
IV-E Domain Adaptation Results
To rigorously assess the efficacy of Strada-LLM compared to prompt-tuning techniques, particularly in few-shot learning scenarios, we conduct a comprehensive evaluation. Our experimental design involves fine-tuning Strada-LLM for a duration equivalent to that used in the FlashST approach, as proposed in [14] on simple universal prompt tuning. The results of this comparative analysis are meticulously documented in Table V, which presents a detailed breakdown of performance metrics for both the Brussels and PeMS07(M) datasets. Upon careful examination of these results, it becomes evident that our Strada-LLM model demonstrates superior performance across the majority of the evaluation metrics when compared to the FlashST method. This improvement in performance serves as a strong indicator of the inherent advantages of our pre-trained, graph-aware LLM in few-shot learning scenarios. The observed superiority can be attributed to Strada-LLM’s ability to effectively leverage its pre-existing knowledge of graph structures, allowing it to adapt more efficiently to new domain with limited training examples. These findings not only validate the effectiveness of our approach but also underscore the potential of graph-aware LLM compared to prompt-based forecasters.
To further validate the adaptability of Strada-LLM, we compare the performance of our proposed method with the most relevant transfer learning approaches from the literature, including:
  • Full fine-tuning: All of the trainable parameters of the model will be fine-tuned.
  • $Top^{k}$ fine-tuning [13]: Just fine-tune the last $k$ layers.
  • LoRA [13]: Using low-rank matrices to fine-tune the model.
There are some other methods like instruction tuning [68], which here are not applicable because of the used modality. Moreover, we intertwine LoRA [13] with our proposed model in order to be adapted to other datasets. We experiment with varying sample percentages of the target dataset, in this case PEMS07(M), to identify the saturation point in the few-shot learning process. The results are shown in Figure 4. The plot shows that training on at least 55% of the target data in a few-shot manner is sufficient to achieve good performance. This makes Strada-LLM a versatile LLM for urban data forecasting. Based on the superior performance on the Brussels city dataset, presented in Table V, one can conclude that Strada-LLM is robust when applied to a new city with a different data distribution.
In few-shot scenarios, the convergence plateau is particularly crucial because the model is learning from a limited amount of data, making it prone to overfitting or failing to generalize beyond the training set. To demonstrate the experiment, Figure 5 can verify our effectiveness after 35 epochs. As shown in Table III, the few-shot learning time for both models are identical. Additionally, both LLMs outperform Step (black line in Figure 5) confirming the superiority of LLMs compared to mono-dataset models.
Figure 4: Domain adaptation plateau for Strada-LLM by applying different approaches on the PEMS07(M) dataset in the few-shot setting. The LLM models(highlighted in red and blue) consistently outperform the fully-supervised model (pink), which demonstrates the added value of an LLM.
Figure 5: Epoch convergence plateau comparison on PEMS07(M) dataset.
Notably, as shown in Figure 4, LLM models (i.e., the red and blue plots) outperform the fully supervised method (the black horizontal line); consequently, it shows the capability of LLMs compared to the traditional supervised models.
TABLE VI: Domain adaptation comparison for demonstrating selection of effective source dataset
Source $\rightarrow$ Target $Top^{k}$ Fine-tuning LoRA[13] Fully-Finetune Fully-Supervised
[…]
Metr-LA Pems-BAY 3.31 5.76 2.64 4.83 2.01 3.89 1.54 3.53
Pems-BAY Metr-LA 6.12 8.10 5.28 7.06 4.31 6.18 3.08 5.79
Metr-LA Pems04 22.36 28.72 22.12 27.58 20.41 26.43 19.26 25.47
Pems04 Metr-LA 5.20 6.56 4.83 6.27 4.01 6.13 3.08 5.79
Pems-BAY Pems04 25.76 30.11 24.71 28.96 23.96 29.83 19.26 25.47
Pems04 Pems-BAY 2.73 5.02 1.96 4.15 1.67 3.67 1.54 3.53
IV-F Perturbation Analysis and Robustness
To evaluate obustness of the Strada-LLM model against noise, we conduct perturbation analysis experiments. We added two types of common random noise to the validation data during the experiment. Random noise obeys the Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ where $\sigma\in(0.2,0.4,0.8,1,2)$. Then, we normalize the values of the noise matrices to be between 0 and 1. The results are shown in Figure 6, where the horizontal axis represents $\sigma$, the vertical axis represents the error, and different colors indicate different evaluation metrics. This benchmark is evaluated for the dataset PEMS04 but other datasets follow the same pattern. According to this experiment, our model is robust to Gaussian noise.
Figure 6: Evaluation metrics for the perturbation analysis. Different $\sigma$ values were examined to test the model robustness.
IV-G Ablation Study
IV-G1 Domain Adaptation
In this ablation study, we examine the domain adaptation capabilities of Strada-LLM by evaluating its performance when different datasets are used as source and target domains. Specifically, we trained Strada-LLM on one dataset (e.g., PeMS-04) and adapted to another (e.g., METR-LA) to assess its generalization across different types of traffic patterns. Table VI presents the results for the domain adaptation process when pre-trained on a source dataset and adapted to the corresponding target dataset. The results presented in Table VI indicate that there is a performance drop when the model is applied to a different target domain. This table can also address selecting efficient pre-training datasets to gain transferability advance for the targeted LLM. Take Pems04 as an example, selecting Metr-LA as the pertaining dataset for LLM ends up the 11.39% superior performance in the adaptation phase compared to Pems-Bay, because of the inherent closer distribution. Therefore, considering Pems-bay and Metr-LA as the pertaining datasets will have more distribution coverage compared to Metr-LA and Pems04.
Figure 7: Latent space embedding of the last layer (before the distribution head) visualized in a 2D space.
IV-G2 Model Interpretation
To better understand Strada-LLM, we visualize the latent space of the LLM representation for different datasets. By doing so, the disparity of dataset distributions and the reason for the difficulty of adapting to PEMS04 are revealed. The visualization is shown in Figure 7 which is obtained with T-SNE [69] in 2D dimensions. Our variational posterior distribution $P(\theta\mid X)$ can truly depict the underlying distribution.
IV-G3 LoRA
@@
We evaluate the performance of our LoRA adoption with different widths. To do so, we try various values for $R\in(2,4,8)$ and the results are demonstrated in Table VII. As can be seen, the lower error comes with the cost of involving more parameters. For real-time applications, the lower bandwidths are more applicable.
TABLE VII: LoRA rank benchmark results on PeMS04 dataset
[…]
$2$ 24.76 29.37 12.80
$4$ 22.12 27.58 12.09
$8$ 21.87 26.62 11.8
IV-G4 LLM Cross-compatibility
We evaluate our model’s flexibility by deploying it across multiple LLM architectures. Specifically, we utilized Mistral and Llama 2 as base models to assess the adaptability of our proposed method. The results, shown in Table VIII, indicate that the framework is adaptable to different LLM backbones without losing too much performance
TABLE VIII: Ablation study on LLM backbone adaptability for PEMS07(M)
[…]
Llama2 [70] 2.81 5.63 6.81
Mistral [38] 2.60 4.86 6.51
V Conclusion
This research proposes a probabilistic-based LLM for traffic forecasting called Strada-LLM, which injects the graph structure implicitly into the input tokens. We utilize a subgraph extraction procedure to inject the graph structure into the transformer token inputs. On the one hand, we adopt a probabilistic transformer to predict the traffic data by sampling from the underlying distribution. On the other hand, we utilize a low-rank method for the transfer learning task. Consequently, we showcase its success in adapting to datasets with significant differences compared to the datasets used for pre-training. Finally, the model is robust to noise, interpretable, and scalable. Strada-LLM is compliant to a variety of different known LLM backbones, which make its architecture suitable to process spatio-temporal graphs. In summary, the Strada-LLM model successfully captures the spatial and temporal features from traffic data so that they can be applied to other spatio-temporal tasks. A potential avenue for future research could involve evaluating the expressiveness of the distribution head.
Acknowledgment
This work is funded by Innoviris within the research project TORRES. The authors thank Leandro Di Bella, Weijiang Xiong, and Bert Van Hauwermeiren for comments that greatly improved the manuscript.

### v3

@@ A.1. Baselines
  • GMAN  (Zheng et al., 2020): The gating fusion mechanism redefines the spatio-temporal attention block.
  • MTGNN  (Wu et al., 2020): Alternating use of graph convolution and temporal convolution modules.
@@
  • GTS  (Shang et al., 2021): The underlying graph structure is learned among multiple time series, and the corresponding time series is simultaneously predicted with DCRNN (Li et al., 2018).
@@
  • STEP  (Shao et al., 2022): Adopts the graph wavenet model (Wu et al., 2019), integrated into the transformer backbone and pre-training scheme.
@@
  • FlashST (Li et al., 2024b), UniST (Yuan et al., 2024): The data distribution shift is learned by leveraging prompt-tuning.
  • Lag-Llama (Rasul et al., 2024): They used known LLM backbones for unified probabilistic time-series forecasting.
@@
To compare with the baselines, we adopt MAE, RMSE, and MAPE prediction error (Li et al., 2024b).
A.2. Proof of Equation LABEL:kl_formula
Let $f_{d}({\bf x}\mid\boldsymbol{\mu},\boldsymbol{\Sigma},\nu)$ be a multivariate $t$, of dimension $d$, with location vector $\boldsymbol{\mu}$, scale matrix $\boldsymbol{\Sigma}$ and $\nu$ degrees of freedom. The aim is to define an objective prior for the parameter $\nu$. For simplicity in the notation, we will write $f_{d,\nu}=f_{d}({\bf x}\mid\boldsymbol{\mu},\boldsymbol{\Sigma},\nu)$, for $\nu=1,\ldots,\nu_{\max-1}$, and $f_{d,\nu_{\max}}=N_{d}({\bf x}\mid\boldsymbol{\mu},\boldsymbol{\Sigma})$, with
$N_{d}({\bf x}\mid\boldsymbol{\mu},\boldsymbol{\Sigma})=\frac{1}{\sqrt{(2\pi)^{d}|\boldsymbol{\Sigma}|}}\exp\left\{-\frac{1}{2}({\bf x}-\boldsymbol{\mu})^{\top}\boldsymbol{\Sigma}^{-1}({\bf x}-\boldsymbol{\mu})\right\},$
where in this case $\boldsymbol{\mu}$ is the vector of means and $\boldsymbol{\Sigma}$ is the covariance matrix. The prior for $\nu$ here discussed depends on the Kullback–Leibler divergence between two multivariate densities. In particular, for $\nu=1,\ldots,\nu_{\max-1}$, the prior is based on the Kullback–Leibler divergence between two multivariate $t$ densities, which differ only in the number of degrees of freedom. The divergence between two $d$-variate $t$ densities, $f_{d,\nu}$ and $f_{d,\nu^{\prime}}$, is given by
$D_{KL}(f_{d}(\cdot\mid\boldsymbol{\mu},\boldsymbol{\Sigma},\nu)\mid\mid f_{d,}(\cdot\mid\boldsymbol{\mu},\boldsymbol{\Sigma},\nu^{\prime}))$
$=$ $D_{KL}(f_{d}(\cdot\mid{\bf 0},{\bf I},\nu)\mid\mid f_{d}(\cdot\mid{\bf 0},{\bf I},\nu^{\prime}))$
$=$ $\int_{{\mathbb{R}}^{n}}f_{d}({\bf x}\mid{\bf 0},{\bf I},\nu)\log\dfrac{f_{d}({\bf x}\mid{\bf 0},{\bf I},\nu)}{f_{d}({\bf x}\mid{\bf 0},{\bf I},\nu^{\prime})}\,d{\bf x}$
$=$ $\int_{{\mathbb{R}}^{n}}K(d,\nu)\left(1+\dfrac{{\bf x}^{\top}{\bf x}}{\nu}\right)^{-\frac{\nu+d}{2}}\log\dfrac{K(d,\nu)\left(1+\dfrac{{\bf x}^{\top}{\bf x}}{\nu}\right)^{-\frac{\nu+d}{2}}}{K(d,\nu^{\prime})\left(1+\dfrac{{\bf x}^{\top}{\bf x}}{\nu^{\prime}}\right)^{-\frac{\nu^{\prime}+d}{2}}}\,d{\bf x}$
(8) $=$ $\log\dfrac{K(d,\nu)}{K(d,\nu^{\prime})}-\dfrac{\nu+d}{2}{\mathbb{E}}_{d,\nu}\left[\log\left(1+\dfrac{{\bf x}^{\top}{\bf x}}{\nu}\right)\right]+$
$\dfrac{\nu^{\prime}+d}{2}{\mathbb{E}}_{d,\nu}\left[\log\left(1+\dfrac{{\bf x}^{\top}{\bf x}}{\nu^{\prime}}\right)\right],$
where
$K(d,\nu)=\dfrac{\Gamma\left(\dfrac{\nu+d}{2}\right)}{\Gamma\left(\dfrac{\nu}{2}\right)\sqrt{(\pi\nu)^{d}}},$
and ${\mathbb{E}}_{d,\nu}$ represents the expected value with respect to $f_{d}(\cdot\mid{\bf 0},{\bf I},\nu)$. More specifically,
${\mathbb{E}}_{d,\nu}\left[\log\left(1+\dfrac{{\bf x}^{\top}{\bf x}}{\nu}\right)\right]=\Psi\left(\dfrac{\nu+d}{2}\right)-\Psi\left(\dfrac{\nu}{2}\right),$
where $\Psi$ is the digamma function. For the second expectation in (A.2), we use Lemma 2 in (Zografos, 1999) to obtain the following expression after a change of variables in terms of spherical coordinates
$\mathbb{E}_{d,\nu}\left[\log\left(1+\frac{\mathbf{x}^{\top}\mathbf{x}}{\nu^{\prime}}\right)\right]$ $=$ $K(d,\nu)\frac{\pi^{\frac{d}{2}}}{\Gamma\left(\frac{d}{2}\right)}\times$
$\int_{0}^{\infty}\left(1+\frac{t}{\nu}\right)^{-\frac{\nu+d}{2}}t^{\frac{d}{2}-1}\log\left(1+\frac{t}{\nu^{\prime}}\right)dt$
which only requires one-dimensional numerical integration, regardless of the dimension $d$.
Figure 7. The visualization of $k$-hop multivariate zero-shot prediction for Pems07(M) dataset for long-term horizon
(a) Node 6
(b) Node 13
(c) Node 48
Figure 8. Strada-LLM traffic prediction in the capturing volatility scenarios for three (not-connected) different nodes exist in Pems07(M) dataset.
A.3. Visualizations
To better understand and assess our model visually, we present additional illustrations in this section. The visualization of long-term forecasting in zero-shot setting for Pems07(M) dataset is shown in Figure 7. Figure 7 is an instance of multivariate probabilistic forecasting in which the dark blue nodes are part of the k-hop extracted subgraph. A critical aspect of time-series prediction algorithms is their ability to capture volatility importance - the varying degrees of uncertainty and rapid changes in the data over time. By incorporating volatility measures, these algorithms can better understand periods of high traffic turbulence versus relative stability, leading to more nuanced and accurate predictions. Strada-LLM is also able to capture volatility existing in Pems07(M) dataset, which is shown in Figure  8.
A.4. Gradient Analysis
As mentioned in section  3.3, the low-rank adaptation has not been utilised in the Graph Embedding module, which causes model performance degradation. To explain the reason, the norm and variance magnitude of the gradients of two MLP layers inside the module are plotted in Figure 9. As shown in Figure 9(b), according to the heterogeneous gradient backpropagation of these layers, e.g., one MLP’s variance is 10 times the other MLP layer, and the adoption of the same low-rank adaptation causes performance degradation. Proposing an efficient low-rank adaptation in this use case will remain an ongoing research direction.
(a) Comparison of $l_{2}$ gradient norms between the first layer (left) and the second layer (right) across training iterations.
(b) Gradient $\sigma^{2}$ variance comparison between LaplacianPE1 (left) and LaplacianPE2 (right) layers during training.
Figure 9. Training dynamics comparison between two graph Laplacian positional encoding layers: gradient norms and variances over iterations.
A.5. Training Strategy
A.5.1. Augmentation Techniques
We adopted the time series augmentation techniques FreqMixFreq\-Mix(Chen et al., 2023) and FreqMaskFreq\-Mask (Chen et al., 2023) to prevent over-fitting. Our training framework comprises two distinct stages: pre-training and fine-tuning. During pre-training, we leverage a diverse combination of datasets: $D_{base}$ for general language understanding, $D_{domain}$ for domain-specific knowledge, and $D_{task}$ for task-oriented capabilities. These datasets are processed in a curriculum learning manner, starting with $D_{base}$ to establish foundational knowledge, followed by $D_{domain}$ to inject domain expertise, and finally $D_{task}$ to enhance task-specific performance. To maintain balanced exposure across datasets while optimizing computational efficiency, we employ a dynamic batch construction strategy where each training batch $B_{t}$ consists of samples drawn from multiple datasets with proportional representation: $\alpha_{1}$% from $D_{base}$, $\alpha_{2}$% from $D_{domain}$, and $\alpha_{3}$% from $D_{task}$, where $\sum_{i}\alpha_{i}=1$. This balanced approach ensures the model develops robust general capabilities while maintaining strong performance on domain-specific tasks. In the fine-tuning stage, we exclusively focus on $D_{task}$ with smaller learning rates and specialized optimization techniques to prevent catastrophic forgetting of pre-trained knowledge while adapting to target objectives.
A.5.2. Experimental Setup
The hyperparameters of the Strada-LLM model mainly include: learning rate, batch size, training epochs, the number of layers, context length, embedding dimension per head, number of heads, rope scaling, and number of parallel samples for greedy decoding. We adopt Adam as the optimizer (Kingma, 2014). In the experiment, we follow the weight decay 1e81e-8 and set the learning rate to 0.001 (Kingma, 2014), the batch size to 32, and the training epochs to 150 for each dataset. We set $k=3$ in the sub-graph extractor. Furthermore, based on the fact that Strada-LLM is a probabilistic model, we take samples and compute medians to compare with point-wise models. The number of samples is set to $100$. In terms of complexity, our model contains 16 million parameters. The experiments are conducted on 2 Nvidia RTX 4090 GPUs.
Table 13. LoRA rank benchmark results on PeMS04 dataset
@@ A.5.2. Experimental Setup
22 24.76 29.37 12.80
44 22.12 27.58 12.09
88 21.87 26.62 11.8
A.6. Rank Analysis
@@
We evaluate the performance of our LoRA adoption with different widths. To do so, we try various values for $r\in(2,4,8)$ and the results are demonstrated in Table 13. As can be seen, the lower error comes with the cost of involving more parameters. For real-time applications, the lower bandwidths are more applicable.
A.7. Perturbation Analysis and Robustness
To evaluate the robustness of the Strada-LLM model against noise, we conduct perturbation analysis experiments. We added two types of common random noise to the validation data during the experiment. Random noise obeys the Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ where $\sigma\in(0.2,0.4,0.8,1,2)$. Then, we normalize the values of the noise matrices to be between 0 and 1. The results are shown in Figure 10, where the horizontal axis represents $\sigma$, the vertical axis represents the error, and different colors indicate different evaluation metrics. This benchmark is evaluated for the dataset PeMS04 but other datasets follow the same pattern. According to this experiment, our model is robust to Gaussian noise.
Figure 10. Evaluation metrics for the perturbation analysis. Different $\sigma$ values were examined to test the model’s robustness.
Figure 11. Strada-LLM scalability analysis compared to existing LLMs.
A.8. Scalability Comparison
Our empirical analysis reveals a nuanced relationship between model size and performance, characterized by diminishing returns beyond certain parameter thresholds. As illustrated in Figure 11, our model performs superior on the similar parameter capacity on the Crowd dataset.